# DESI Dark Energy Time Evolution is Recovered by Cosmologically Coupled Black Holes

### v1

@@ 4 Discussion
Although the amount of baryonic density consumed is technically viable, the astrophysical processes directing it towards BHs are unclear. Assuming a Chabrier distribution of stellar masses at birth $\mathrm{d}N/\mathrm{d}m\propto m^{-2.3}$, the fraction of stellar mass in stars large enough to form BHs ($m\gtrsim 20M_{\odot}$) is $f=0.4$. If this fraction of stellar mass becomes BH mass, the value of $\Xi$ implies an additional growth factor of approximately 12.5$\sim$ 12.5$\times$ 12.5 × is required. We consider some channels that could contribute to this factor. The stellar Initial Mass Function (IMF) at $z\gtrsim 1$ is an active area of investigation [e.g. 96, 97]. If the IMF at $z\gtrsim 1$ were ‘top-heavy,’ as suggested by some studies [98, 99, 100, 101], the required factor decreases. For example, an IMF $\propto m^{-2.1}$ [e.g. 102] gives $f=0.74$, reducing the factor to 6.7×6.7$\times$6.7 ×. Post-collapse accretion could account for part of $\Xi$, though estimating its extent is challenging. Accretion onto isolated BHs is believed to be inefficient. Because $\Xi$ describes average behavior, a scenario where the high-end of the BH mass function accretes by a factor of $\sim 10^{2}\text{\textendash}10^{3}\times$, while the low end evolves passively, is conceivable. Gravitational processes can also produce mass. This physics is independent from $\Xi$ because Eq. 2.3 has no knowledge of local gravitational processes. The Press and Teukolsky process [103, 104] can convert BH spin into mass, provided radiated energy is absorbed by further infalling material during core collapse. Local strong gravitational effects would effectively amplify the term on the right-hand side of Eq. 2.3, without altering the depletion described by Eq. 2.1. We note that such effects may involve mediating fields, such as scalar fields [105, 106].
The “cosmic noon” scenario we consider here differs from the previously considered “cosmic dawn” scenarios, where high-$z$ BH production leveraged the $\sim 1/a^{3}$ factor in Eq. 2.9 to achieve $\Omega_{\Lambda}\rho_{\mathrm{cr}}$ from a small fraction of baryons [107, 61]. We note that any contribution from Population III stellar collapse and direct-collapse BHs at $10\lesssim z\lesssim 25$ would behave, at background order, as an effective cosmological constant. A scenario where $0.5\Omega_{\Lambda}$ is established at high-$z$ would, in principle, reduce $\Xi$ by an additional factor of two. Despite the agreement in shape, it is evident that further work is required to understand and estimate baryon consumption for a “cosmic noon” scenario.
In summary, DESI data suggest time-evolution in the dark energy density at $2.5\operatorname{-}3.9\sigma$ confidence when combining its Baryon Acoustic Oscillation measurements with Cosmic Microwave Background and various supernovae datasets. The time-evolution is characterized by a two-parameter dark energy equation of state $w=w_{0}+w_{a}(1-a)$, where $a$ is the scale factor. Our analysis shows that the DESI most-likely values for $w_{0}$ and $w_{a}$ are recovered when a $w_{0}w_{a}$ model is fit to dark energy produced by baryon conversion in cosmologically coupled black holes. Baryon conversion is constrained solely by measurements of the comoving star formation rate density, which are independent of the DESI measurements and previous observational studies of cosmologically coupled black hole solutions. As DESI gathers more data, the increased statistical precision will further constrain $w_{0}$ and $w_{a}$, potentially enabling DESI to measure the time evolution of dark energy independently, without relying on external Type Ia Supernova data.
@@ Acknowledgments
KC thanks the Sakai Group at Yamaguchi University for their hospitality during the preparation of this work. KC further thanks Ryo Saito (Yamaguchi) for comments on $w_{0}w_{a}$ priors, Joel Weiner (Hawai‘i) for discussion, and Tom Browder (Hawai‘i) for comments on neutrinos. GT acknowledges support through DoE Award DE-SC009193. RAW acknowledges support from NASA JWST Interdisciplinary Scientist grants NAG5-12460, NNX14AN10G and 80NSSC18K0200 from GSFC. The work of NF is supported by DOE grant DOE-SC0010008. NF thanks the Aspen Center for Physics, which is supported by the National Science Foundation grant PHY-2210452, and the Sloan Foundation for its partial support.

### v2

@@ 5 Discussion
To place the measured value of $\Xi$ into context, we assume a Chabrier [113] distribution of stellar masses at birth $\mathrm{d}N/\mathrm{d}m\propto m^{-2.3}$. Then the fraction of stellar mass in stars large enough to form BHs ($m\gtrsim 20M_{\odot}$) is $f=0.4$. A $\Xi=1.4$ implies approximately 3.45×3.45$\times$3.45 × additional baryonic consumption, which is reasonable given astrophysical uncertainties. For example, the stellar Initial Mass Function (IMF) at $z\gtrsim 1$ is an active area of investigation [e.g. 114, 115]. If the IMF at $z\gtrsim 1$ were ‘top-heavy,’ as suggested by some studies [116, 117, 118, 119], the required factor decreases. For example, an IMF $\propto m^{-2.1}$ [e.g. 120] gives $f=0.74$, reducing the factor to 1.86×1.86$\times$1.86 ×. Post-collapse accretion could account for part of $\Xi$, though estimating its extent is challenging. Accretion onto isolated BHs is believed to be inefficient. Because $\Xi$ describes average behavior, a scenario where the high-end of the BH mass function accretes by a factor of $\sim 10^{2}\text{\textendash}10^{3}\times$, while the low end evolves passively, is conceivable. Gravitational processes can also produce mass. This physics is independent of $\Xi$ because Eq. 2.4 has no knowledge of local gravitational processes. The Press and Teukolsky process [121, 122] can convert BH spin into mass, provided the radiated energy is absorbed by further infalling material during core collapse. Local strong gravitational effects would effectively amplify the term on the right-hand side of Eq. 2.4, without altering the depletion described by Eq. 2.2. We note that such effects may involve mediating fields, such as scalar fields [123, 124].
We have focused on a “cosmic noon” scenario, but production of DE at the peak of star-formation complements production near “cosmic dawn” $z\sim 20$, as has been previously studied [125, 63, 64, 68]. The impact of Pop III [e.g. 126] and direct-collapse BHs [e.g. 127] is to establish an effective DE “floor.” Production of this DE floor does not significantly deplete the baryon density, because the DE produced per unit baryon density scales as $(1+z)^{3}$. For example, in Planck 2018 cosmologies, $\omega_{\Lambda}\sim 5\omega_{c}/2$. Assume a cosmic dawn burst of production, either direct collapse or Pop III stellar collapse, at $z_{\mathrm{III}}$. Then
$\frac{5\omega_{c}}{2}=\omega_{b}\left(1-s_{\mathrm{III}}\right)\left(1+z_{\mathrm{III}}\right)^{3}.$ (5.2)
Adopting $z_{\mathrm{III}}=15$ then gives a baryon survival of $s_{\mathrm{III}}=99.67\%$, so consumption of $0.4\%$ of baryons is sufficient to produce all required DE. Any DE floor established at high $z$ would further decrease $\Xi$.
In summary, DESI’s first year of observations provide evidence that DE evolves over time using a $w_{0}w_{a}$ parameterization. If DE is produced from the conversion of baryonic matter into cosmologically coupled BHs, the DE density will increase as massive stars form and collapse into BHs, rising from zero toward a plateau as star-formation quenches at $z\lesssim 2$. We have tested this hypothesis by fitting a suitable Friedmann cosmology to the DESI measured Baryon Acoustic Oscillation signals between $2.50<z<$ 2.5. Our calculated evolving DE density agrees with the DESI $w_{0}w_{a}$ models at $<1\sigma$, except at redshifts $z\lesssim 0.2$, where the $w_{0}w_{a}$ parameterization becomes inadequate for our phenomenology. Cosmologically coupled BHs produce $H_{0}=(69.94\pm 0.81)\,{\rm km\,s^{-1}\,Mpc^{-1}}$, in excellent agreement with $H_{0}=(69.59\pm 1.58)\,\,{\rm km\,s^{-1}\,Mpc^{-1}}$ reported very recently by Chicago-Carnegie Hubble Program using Cepheids, Tip of the Red Giant Branch (TRGB), and J-Region Asymptotic Giant Branch stellar distance-ladder calibrations. They provide a physical explanation for this larger central value: the Hubble rate today is higher than that inferred from the Cosmic Microwave Background because the conversion of baryons into DE during star-formation causes an earlier transition from matter to DE dominance. In addition to reducing tensions between early and late-time observations, the scenario we have investigated provides an astrophysically motivated explanation for a known $\sim 30\%$ deficit in the present-day baryon density, relative to BBN expectations. Furthermore, the conversion of baryons into DE decreases the present-day matter density of Friedmann models, allowing the sum of neutrino masses to increase away from zero. No other current DE model is both astrophysically motivated and capable of reconciling early and late universe measurements of the expansion rate.
@@ Acknowledgments
The authors thank the anonymous referee for very useful comments that resulted in substantial improvement of the manuscript. We further thank Arnaud de Mattia and Seshadri Nadathur (DESI) for discussion concerning regeneration of DESI results, Thomas Tram (Norway) for guidance with massive relics in CLASS, Mark Dickinson (NOIRLab) for guidance with cosmological corrections to SFRDs, and Christopher Cain (ASU) for input on reionization. KC thanks the Sakai Group at Yamaguchi University for their hospitality during the preparation of this work, Ryo Saito (Yamaguchi) for comments on $w_{0}w_{a}$ priors, Joel Weiner (Hawai‘i) for discussion, and Tom Browder (Hawai‘i) for comments on neutrinos. GT acknowledges support through DoE Award DE-SC009193. RAW acknowledges support from NASA JWST Interdisciplinary Scientist grants NAG5-12460, NNX14AN10G and 80NSSC18K0200 from GSFC. The work of NF is supported by DOE grant DOE-SC0010008. NF thanks the Aspen Center for Physics, which is supported by the National Science Foundation grant PHY-2210452, and the Sloan Foundation for its partial support.